OpenStudy (anonymous):

Find the zeros: x^2+3x-10=0. I get the answer x=-5. Is this correct?

OpenStudy (anonymous):

(-5)^2+3(-5)-10 = 25 - 15 - 10 = 0. correct!

OpenStudy (anonymous):

that's not the only 0, however

OpenStudy (anonymous):

What is the other 0?

OpenStudy (anonymous):

Is the other answer 2?

OpenStudy (anonymous):

yes